OpenStudy (anonymous):

In a chemical reaction. How do we know which of the two substance is IN EXCESS?

OpenStudy (callisto):

determine the number of mole of each reactant, the smallest one is the limiting agent, while the other reactants are in excess

OpenStudy (callisto):

eg, 10g of MgO reacts with 100cm^3 of 0.1M HCl MgO + 2HCl -> MgCl2 +H2O number of mole of MgO used = 10/(24.3+16.0) = 0.025720mole number of mole of HCl used = 100/1000 x 0.1 = 0.01 mole since 1mole of MgO required 2 mole of HCl for complete reaction. 0.025720mole of MgO needs 2x0.025720mole = 0.05144 mole of HCl for complete reaction However we have only 0.01mole of HCl for the reaction, which is not enough. Therefore, HCl is the limiting agent. On the contrary, MgO is in excess.

OpenStudy (anonymous):

you know that nitrogen and hydrogen reacts in ratio 1 to 3. suppose nitogen given is 1 mole and hydrogen given is 2 mole. clearly hydrogen is less....so to find limitung reagent , you have to consider the stoicometric ratios and use it to find which is in excess, is it clear now?